OpenStudy (anonymous):

Solve. 3x2 + 6x = 8?

OpenStudy (anonymous):

is that 3x squared?

OpenStudy (anonymous):

The answers are in quadratic formula form, so the quadratic formula needs to be used to solve this. I dont want to cheat, I just need help...Im confused. What I did: subtracted 8 from both sides to get all terms on one side. plugged the numbers into the q-formula, as a=3, b=6, and c=8. I worked it down to x=-6 +/- radical 132 /6...and now Im stuck. I dont know if I did anything wrong ir what to do next...Please help?

OpenStudy (anonymous):

Yes, it is 3x squared.

OpenStudy (lgbasallote):

you subtracted 8 from both sides...that means c = -8 not positive

OpenStudy (anonymous):

Oh yeah, I did the problem that way...I just wrote it wrong here. In the radical negative 8 is multiplied by 3 and negative 4, which get's 132 in the radical.

OpenStudy (anonymous):

Write the equation as \[ 3 x^2 + 6 x - 8 == 0 \] Use the quadratic formula to find \[ x_1=\frac{1}{3} \left(-3-\sqrt{33}\right)\\ x_2=\frac{1}{3} \left(-3 +\sqrt{33}\right) \]

OpenStudy (anonymous):

When\[ b=2 b'\] , the quadratic formul simplifies to \[ a=\frac {-b' \pm \sqrt{b'^2 - ac} } {a} \]

OpenStudy (anonymous):

the last line a should be x